# Correctness Assessment of Code Generated by Large Language Models Using Internal Representations

### v1

@@ 5.1.3 Error analysis
TABLE IV: Openia’s performance on different errors
@@
Table IV provides insights into Openia’s ability to predict correct and incorrect code by analyzing different error categories. In this table, the Correct row represents Openia’s performance in accurately predicting correct code. Meanwhile, the remaining rows are the results of Openia in detecting incorrect code, categorized by error types.
@@
Overall, Openia demonstrates higher accuracy in detecting incorrect code compared to the correct one. Specifically, Openia’s F1-Score for identifying correct code is 0.75, whereas its performance in detecting incorrect code reaches 0.82. This is reasonable because identifying incorrectness often requires just a single signal/pattern indicating an error, making it easier to classify a code snippet as incorrect. In contrast, determining whether a program is correct requires assessing multiple factors such as logic, structure, and functionality, which makes it a more challenging task.
@@
Among the error types, Openia exhibits the lowest performance in detecting incorrect code caused by index out of range errors. Specifically, its F1-Score for identifying incorrect code of this error is 0.73, which is 12% lower than its average performance of detecting incorrect code. Unlike static errors, such as undefined variables or type errors, that are inherently present in the code and can be directly captured in the LLM’s internal states, index out-of-range errors are not always explicitly encoded in the internal representations of the model. This is because such errors often depend on dynamic factors like runtime conditions or specific input values. Therefore, to effectively detect incorrect code of this error, it could require a deeper understanding of program execution and data dependencies.
@@ 5.2.1 Layer selection analysis
Fig. 3 shows the performance of Openia in correctness assessment when using hidden states from different layers of three code LLMs: MagicCoder 7B, CodeLlama 7B, and DeepSeek Coder 6.7B. In general, the representations from the middle layers of the models enable Openia to achieve optimal performance for both independent-unit code generation and repo-level code generation tasks. For independent-unit code generation (Fig. 3(a)), Openia obtains its best performance by leveraging the representations from layer 28 of MagicCoder, layer 16 of CodeLlama, and layer 20 of DeepSeek. Similarly, for repo-level code generation (Fig. 3(b)), Openia attains its highest F1-Score using the representations from layer 28 for both MagicCoder and DeepSeek. In the case of CodeLlama, while Openia’s peak performance is achieved with the representations from the last layer (layer 32), the middle layers (layers 16-28) still offer consistently strong results. These findings suggest that the middle layers in the LLMs typically capture a balance between the local context focused by the shallow layers and the global context encoded by the deep layers, making them optimal for assessing code correctness.
@@ 5.2.2 Token selection analysis
Fig. 4 shows the performance of Openia when using token representations from different layers of three models, MagicCoder 7B, CodeLlama 7B, and DeepSeek Coder 6.7B. In all three models, the first layer corresponds to layer 1, the middle layer is layer 16, and the last layer is layer 32. For each layer, we extract and evaluate the performance of Openia with the representations of tokens at four specific positions: first token and last token, which represent the starting and ending points of the entire answer generated by the LLM, and first code token and last code token, which correspond to the starting and ending points of the code segment within the LLM’s response.
In general, the deeper layer, the more stable Openia’s performance is across different token representations; conversely, the shallower layer, the more sensitive Openia becomes to variations in token selection. Specifically, when using the representations from the last layers, Openia consistently achieves the highest F1-Score with the representations of the last code token. Furthermore, its performance across different token representations from the last layers remains relatively stable. In contrast, when using representations from the first layer, there is a noticeable performance gap between different tokens. For example, in assessing the correctness of code generated by MagicCoder for independent-unit tasks, (Fig. 4(a)), Openia achieves an F1-Score of 0.7 with the last token representations from the first layer, which is 30% higher than its performance with the first code token representations.
@@
However, the impact of token selection on Openia’s performance varies depending on the model and layer. For instance, MagicCoder exhibits less sensitivity to token selection, while CodeLlama is highly affected. For example, with the representations from the last layer of MagicCoder, Openia’s F1-Score in predicting the code correctness for independent-unit code generation ranges from 0.74 to 0.78. However, these figures for CodeLlama vary more widely, from 0.66 to 0.75. This highlights the importance of carefully selecting suitable tokens to extract representations to obtain the best performance of Openia. Similar to layer selection, token selection should be dynamically adapted to align with the specific architecture and characteristics of the model.
@@ 5.3.1 Programming language analysis
TABLE V: Correctness assessment performance in F1-Score of the approaches across different languages
@@
As seen in Table V, Openia consistently outperforms the traditional classification-based approaches that encode the final generated code of the LLMs using pre-trained language models such as CodeBERT and CodeT5+. Across all the studied programming languages, Openia achieves an average F1-Score of 0.75, which is 12% higher than CodeT5+ and 23% higher than CodeBERT. Notably, for Python, Openia surpasses CodeT5+ and CodeBERT by 35% and 81%, respectively. These results indicate that the internal representations of the LLMs can effectively capture not only the code semantics but also the fundamental programming principles encoded in the code generation process of the model. Levering the internal representations of the LLMs enables Openia to generalize its learned knowledge and better detect code correctness across different languages.
All the approaches achieve their highest performance when the target language of the test set is TypeScript (TS) and their lowest performance when the target language is Python. For example, Openia’s F1-Score for predicting the correctness of Python code is 0.67, which is 12% lower than its average score. Meanwhile, for TS code, Openia achieves an impressive F1-score of 0.88, which is 17% higher than its average score. A similar trend is observed with CodeBERT and CodeT5+. This is due to Python’s inherent characteristics, such as dynamic typing, indentation-based syntax, and pythonic idioms, which differ significantly from the other languages. These features introduce ambiguity and variability, making it harder for Openia to generalize its learned knowledge to Python code. In contrast, TS shares many common characteristics with other languages, such as supporting both static and dynamic typing, verbose and explicit code, and curly braces-based syntax. As a result, Openia can effectively generalize its learned knowledge from other languages to accurately predict the correctness of TS code.
@@ 5.3.2 Task difficulty
In this experiment, we investigate how the difficulty levels of the input tasks affect Openia’s performance. First, we prompt the LLM to classify each task in the HumanEval dataset into one of five difficulty levels: Very Easy, Easy, Medium, Hard, and Very Hard. The prompt used and the assigned difficulty levels for each task are available on our website [36]. Then, we analyze the performance of Openia in predicting the correctness of the generated code for tasks within each difficulty group.
As shown in Fig. 5, Openia obtains higher performance on both easy and hard tasks, while its performance declines for medium-level tasks. Specifically, for easy tasks, Openia correctly predicts the correctness of generated code with an F1-score of 0.75, which is 25% higher than its performance on medium-level tasks. These findings suggest that the internal representations of the LLM can capture its confidence about its responses. For instance, the LLM tends to be more confident when handling easy tasks and less confident with hard ones. This confidence signal can effectively help to assess code correctness, as reflected in Openia’s high performance on these tasks. However, for medium-level tasks, the confidence signals appear to be less distinct, leading to a decrease in Openia’s performance.
@@ 5.4 Time complexity
The time complexity of Openia is primarily influenced by the extraction of internal representations from the studied code LLMs, training the lightweight probing classifier, and the inference process.
Extracting the internal representations from the studied code LLMs was conducted on an NVIDIA A6000 GPU with 48GB of VRAM. This process required approximately 86 hours for the entire dataset of 24,890 code units generated by the LLMs. Meanwhile, all the other experiments were conducted on a single NVIDIA P100 GPU with 16GB of VRAM. Openia is highly efficient in terms of both training and inference. With a lightweight probing classifier consisting of two hidden layers (128 and 64 neurons), the training process takes approximately 1 minute, as there is no embedding computation overhead. In contrast, classifiers built on pre-trained models such as CodeBERT and CodeT5+ require significantly more time, taking about 83 minutes for training. The inference process of Openia requires only 0.6 milliseconds per generated code unit. This demonstrates the feasibility of Openia for real-time, in-process assessment of code correctness during generation by LLMs, offering substantial advantages in practical deployment scenarios.
@@ 5.5 Threats to Validity
Internal Validity. Threats to internal validity include the selection of hyperparameters for the baseline methods and Openia. To mitigate this threat, we systematically explored different hyperparameter configurations for Openia and evaluated its performance under varied conditions (as detailed in Section 5). For baseline methods, we reused implementations and settings from the original papers, such as those for CodeBERT and CodeT5+ [40, 42], ensuring fair and consistent comparisons. Additionally, potential threats could arise from the procedures used to analyze the internal signals of LLMs and correlate them with code correctness. To address this, we implemented robust and well-documented methods to extract and interpret the intermediate states of LLMs during code generation. These methods were validated on diverse benchmarks and cross-verified using established practices in the field. To further minimize the risk of errors in implementation, our codebase has undergone extensive internal review, and we have made it publicly available [36].
@@
Construct Validity. Threats to construct validity relate to the appropriateness of our evaluation metrics and procedures. We employed widely recognized metrics, including Accuracy, Precision, Recall, and F1-Score, to measure code correctness, as these are standard benchmarks in evaluating code generation models. Additionally, we designed experiments to include both standalone and repository-level code generation tasks, ensuring comprehensive evaluations. However, our evaluation primarily relies on automated metrics and does not yet incorporate human assessments. While automated metrics provide objectivity and scalability, human evaluation could offer deeper insights into the practical utility of Openia in real-world coding scenarios. We plan to include such assessments in future work, engaging experienced developers to evaluate the generated code based on factors like functionality, readability, and adherence to programming standards. Another potential limitation is that our controlled experimental setup may not fully reflect real-world software development environments. To address this, we evaluated Openia on multiple representative benchmarks, covering diverse programming constructs and code generation requirements.
@@
External Validity. Threats to external validity concern the generalizability of our results to different LLMs, programming languages, and application domains. To mitigate this, we selected three widely used and representative open-source code LLMs: DeepSeek-Coder, CodeLlama, and MagicCoder [16, 17, 18]. These models are recognized for their strong performance in diverse code-related tasks, making our findings broadly applicable within the code generation domain. A further potential threat arises from focusing primarily on models with parameters up to 7B due to hardware limitations. While this constraint ensures feasibility within our experimental setup, it may limit the applicability of our findings to larger models. Future work will explore scalability to larger models and extend experiments to a broader range of programming languages and codebases to enhance generalizability.

### v2

@@ 5.1.3 Error Analysis
Table 4: Openia’s performance on different errors
@@
Table 4 presents a detailed evaluation of Openia’s performance across different error categories. To compute F1-Score for each error type, we first categorize the generated code snippets into distinct error types according to their ground truth testing results. For each error category, we measure how accurately Openia identified the corresponding code snippets as incorrect. In this table, the Correct row represents Openia’s performance in accurately recognizing correct code, while the remaining rows reports Openia’s performance in detecting incorrect code in each error type.
@@
Overall, Openia demonstrates higher accuracy in detecting incorrect code compared to the correct one. Specifically, Openia’s F1-Score for identifying correct code is 0.75, whereas its performance in detecting incorrect code reaches 0.82. This is reasonable because identifying incorrectness often requires detecting only a single error-indicating signal/pattern, making it relatively easier to classify a code snippet as incorrect. In contrast, determining correctness is inherently more challenging, as it necessitates verifying multiple aspects, including logic, structure, and functionality, to ensure the absence of errors.
@@
Among the error types, Openia exhibits the lowest performance in detecting incorrect code caused by index out of range errors. Specifically, its F1-Score for identifying incorrect code of this error is 0.73, which is 12% lower than its average performance of detecting incorrect code. Unlike static errors, such as undefined variables or type errors, that are inherently present in the code and can be directly captured in the LLM’s internal states, index out-of-range errors are not always explicitly encoded in the internal representations of the model. This is because such errors often depend on dynamic factors like runtime conditions or specific input values. Therefore, effectively detecting incorrect code of this error could require a deeper understanding of program execution and data dependencies.
Table 4 shows our analysis on various error categories for evaluation purposes. While Openia outputs a binary classification, indicating whether the generated code is correct or not, it does not provide detailed diagnostics such as error types or bug locations. As a result, in cases where the code is predicted to be incorrect, developers still need to rely on complementary analysis techniques, such as static analysis, dynamic testing, or multi-agent frameworks, to identify the root cause and guide repair. Despite this limitation, Openia offers practical benefits in several scenarios. First, it can be integrated into the code generation pipeline to proactively filter promising candidates and discard low-quality ones, thereby reducing the risk of presenting incorrect code to developers and minimizing debugging efforts. Second, Openia can serve as a lightweight examiner to determine whether more computationally expensive diagnostic processes are necessary. For future work, we plan to extend Openia by integrating bug localization and repair phases to enhance its practical usages.
5.1.4 Potential in Improving LLM-gen. Code’s Quality
Table 5: Performance of Code LLMs when applying Correctness-Guided Generation (CG2) using Openia (pass@1)
Without CG2 With CG2
DeepSeek Coder-6.7B Cross-benchmark 0.75 0.77
Cross-task 0.63 0.67
Repo-level 0.26 0.28
Code Llama-7B Cross-benchmark 0.60 0.61
Cross-task 0.51 0.57
Repo-level 0.22 0.24
Magicoder-7B Cross-benchmark 0.74 0.76
Cross-task 0.69 0.73
Repo-level 0.24 0.27
To demonstrate the potential of Openia in proactively detecting and preventing bugs before they reach developers, we introduce a Correctness-Guided Code Generation (CG2) pipeline. For each programming task, a LLM is employed to generate several candidates. Openia is then used to assign the correctness score to each candidate, and the highest-scoring one is selected and presented to developers. We compare this pipeline against the standard generation process using the pass@1 metric [33]. Note that, because Openia extracts internal states during generation and feeds them to a lightweight probing classifier with only two hidden layers, the additional time overhead is marginal (more details in Sec. 5.4). Table 5 shows the performance (in pass@1) of the code LLMs, with and without CG2, across different evaluation settings. As seen, Openia is effective in guiding the selection of high-quality code, as demonstrated by the consistent improvements of the LLMs with CG2 over the LLMs without CG2. By selecting the highest-scoring candidates predicted by Openia, the average pass@1 of the code LLMs across different settings improves by about 7%.
We also evaluate a filtering variant in which any solution labeled “incorrect” by Openia is proactively discarded before developer review. Compared to the standard code generation process, this filtering pipeline improves the average rates of passed tests by 6%–86%. Especially, in Repo-level setting, the improvements are at least 53%.
Overall, these results highlight the potential of incorporating Openia into the code generation pipeline to enhance the quality of the generated code.
@@ 5.2.1 Layer Selection Analysis
Figure 5 shows the performance of Openia in correctness assessment when using hidden states from different layers of three code LLMs: Magicoder-7B, Code Llama-7B, and DeepSeek Coder-6.7B. In general, the representations from the middle layers of the models enable Openia to achieve optimal performance for both independent-unit and repo-level code generation tasks. For independent-unit code generation (Fig. 5(a)), Openia obtains its best performance by leveraging the representations from layer 28 of Magicoder, layer 16 of Code Llama, and layer 20 of DeepSeek Coder. Similarly, for repo-level code generation (Figure 5(b)), Openia attains its highest F1-Score using the representations from layer 28 for both Magicoder and DeepSeek Coder. For Code Llama, while Openia’s peak performance is achieved with the representations from the last layer (layer 32), the middle layers (layers 16-28) still offer consistently strong results. These findings suggest that the middle layers in the LLMs typically capture a balance between the local context focused by the shallow layers and the global context encoded by the deep layers, making them optimal for assessing code correctness.
@@ 5.2.2 Token Selection Analysis
Figure 6 shows the performance of Openia when using representations of different tokens from different layers of three models, Magicoder-7B, Code Llama-7B, and DeepSeek Coder-6.7B. In all three models, the first layer corresponds to layer 1, the middle layer is layer 16, and the last layer is layer 32. For each layer, we extract and evaluate the performance of Openia with the representations of tokens at four specific positions: first token and last token, which represent the starting and ending points of the entire answer generated by the LLMs, and first code token and last code token, which correspond to the starting and ending points of the code segment within the LLMs’ response.
In general, the deeper layer, the more stable Openia’s performance is across the internal representations from different tokens; conversely, the shallower layer, the more sensitive Openia becomes to variations in token selection. Specifically, when using the representations from the last layers, Openia consistently achieves the highest F1-Score with the representations of the last code token. Furthermore, its performance across different token representations from the last layers remains relatively stable. In contrast, when using representations from the first layer, there is a noticeable performance gap between different tokens. For example, in assessing the correctness of code generated by Magicoder for independent-unit tasks, (Fig. 6(a)), Openia achieves an F1-Score of 0.7 with the last token representations from the first layer, which is 30% higher than its performance with the first code token representations.
@@
However, the impact of token selection on Openia’s performance varies depending on the model and layer. For instance, Magicoder exhibits less sensitivity to token selection, while Code Llama is highly affected. For example, with the representations from the last layer of Magicoder, Openia’s F1-score in predicting the code correctness for independent-unit code generation ranges from 0.74 to 0.78. However, these figures for Code Llama vary more widely, 0.66–0.75. This highlights the importance of carefully selecting suitable tokens to extract representations to obtain the best performance of Openia. Similar to layer selection, token selection should be dynamically adapted to align with the specific architecture and characteristics of the model.
@@ 5.3.1 Programming Language Analysis
Table 6: Correctness assessment performance in F1-Score of the approaches across different languages
@@
Table 6 shows the generalization performance of Openia and post-hoc classification-based approaches in assessing the code correctness across different programming languages. This experiment is conducted on the multi-language version of HumanEval benchmark. Specifically, for each task in each programming language, DeepSeek Coder-6.7B is employed to generate 10 candidate solutions. To evaluate cross-language generalization, we select one target programming language for testing, and use the generated code snippets (for post-hoc classifier) or internal states (for Openia) from the remaining languages for training. In addition, the row labeled “MIX” in the table corresponds to a setting where all generated code across all languages is combined and randomly split into training and testing sets using a 9:1 ratio. This setting enables to evaluate performance of approaches on mixed language data.
Openia consistently outperforms the traditional post-hoc methods that rely on CodeBERT or CodeT5+ to encode the final LLM-generated code. Across all the studied programming languages, Openia achieves an average F1-Score of 0.75, which is 12% higher than CodeT5+ and 23% higher than CodeBERT. Notably, for Python, Openia surpasses CodeT5+ and CodeBERT by 35% and 81%, respectively. These results indicate that the internal representations of the LLMs can effectively capture not only the code semantics but also the fundamental programming principles encoded in the code generation process. By levering these representations, Openia can generalize its learned knowledge to better detect code correctness across different languages.
Furthermore, all the approaches achieve their highest performance when the target language of the test set is TypeScript and their lowest performance when the target language is Python. For example, Openia’s F1-Score for predicting the correctness of Python code is 0.67, which is 12% lower than its average score. Meanwhile, for TypeScript code, Openia achieves an impressive F1-score of 0.88, exceeding its average score by 17%. A similar trend is observed with CodeBERT and CodeT5+. This is due to Python’s inherent characteristics, such as dynamic typing, indentation-based syntax, and Pythonic idioms, which significantly differ from the other languages. These features introduce ambiguity and variability, making it harder for Openia and the other approaches to generalize their learned knowledge to Python code. In contrast, TypeScript shares many common characteristics with other languages, such as supporting both static and dynamic typing, verbose and explicit code, and curly braces-based syntax. These shared characteristics allow Openia and the other approaches to generalize their learned knowledge more effectively, resulting in higher performance when predicting the correctness of TypeScript code.
@@ 5.3.2 Task Difficulty
In this experiment, we investigate how the difficulty levels of the input tasks affect Openia’s performance. To categorize task difficulty, we first prompt the studied code LLMs to classify each task in the HumanEval dataset into one of five levels: Very easy, Easy, Medium, Hard, and Very hard. The prompt used and the assigned difficulty levels for each task are available on our website [36]. We then analyze the performance of Openia in predicting the correctness of the generated code for tasks within each difficulty group.
As shown in Fig. 7, Openia obtains higher performance on both easy and hard tasks, while its performance declines for medium-level tasks. Specifically, for easy tasks, Openia correctly predicts the correctness of generated code with an F1-score of 0.75, which is 25% higher than its performance on medium-level tasks. These findings suggest that the internal representations of the code LLMs can capture its confidence about its responses. For instance, the code LLMs tend to be more confident when handling easy tasks and less confident with hard ones. This confidence signal can effectively help to assess code correctness, as reflected in Openia’s high performance on these tasks. However, for medium-level tasks, the confidence signals appear to be less distinct, leading to a decrease in Openia’s performance.
5.3.3 Prompt Variability
As shown in Table 2 (Cross-benchmark), differences in prompt templates between training set (MBPP) and testing set (HumanEval) could result in a decline in Openia’s performance, this is not observed for (output-based) post-hoc methods. For DeepSeek Coder-6.7B, F1-score is 0.79 under the Cross-task setting where the prompt template remains consistent between train and test, while the corresponding figure is 0.67 when the prompt templates in the two sets are different. This is reasonable since LLMs’ internal representations inherently encode prompt-specific characteristics, e.g., wording style, instruction phrasing, and contextual emphasis. Consequently, shifts in prompt format induce subtle changes in these latent signals, which in turn affect the features upon which Openia’s classifier depends.
To mitigate this sensitivity, we plan to augment the training set with diverse paraphrases of each prompt. This approach should stabilize Openia’s internal-feature space and further bolster its robustness across varied prompt formulations. Preliminary experiments with mixed-prompt training show an improvement with accuracy increasing from 0.68 to 0.72, confirming that prompt augmentation can enhance Openia’s robustness to prompt variability.
5.3.4 Training Data Size
To understand how the amount of training data affects Openia’s performance, we conducted a controlled experiment using DeepSeek Coder-6.7B in the cross-task setting. We held the test set fixed and partitioned the remaining training data into five equal folds. We then trained Openia on an incremental sequence of folds, from one fold (20% of the data) up to all five folds (100%).
As shown in Fig. 8, both accuracy and F1-Score rise as the training set increases from one to three folds, climbing from 0.71 (accuracy) and 0.67 (F1-score) at one fold to 0.77 for both metrics at three folds. Beyond three folds, performance gains plateau: four folds yield only marginal improvement, and the full five-fold model performs similarly. This trend indicates that Openia quickly learns the key patterns needed to distinguish correct from incorrect code, achieving near-optimal performance with only 60% of the available training data. Adding more data beyond that point produces diminishing returns, suggesting that Openia is data-efficient and can be effectively trained with a relatively small corpus. In practice, practitioners can achieve robust correctness assessment without requiring large labeled datasets, reducing annotation cost and accelerating deployment in resource-constrained settings.
Figure 8: Openia’s performance on training data size
@@ 5.4 Time Complexity
The time complexity of Openia is primarily determined by three factors: preparing training data and extracting internal representations from the studied code LLMs, training the probing classifier, and performing inference.
Preparing the training data for Openia and the studied post-hoc approaches on an NVIDIA A6000 GPU with 48GB of VRAM took approximately 86 hours for the entire dataset of 24,890 code units generated by the LLMs. The extraction of internal representations from the studied code LLMs was performed during code generation without affecting the process of the code LLMs. All subsequent experiments, including probing classifier training and inference, were conducted on a single NVIDIA P100 GPU with 16GB of VRAM. Openia demonstrates high efficiency in both training and inference due to its lightweight probing classifier, which consists of two hidden layers (128 and 64 neurons). The training process takes approximately 1 minute, as there is no embedding computation overhead. In contrast, classifiers built on pre-trained models, such as CodeBERT and CodeT5+, require significantly more time, taking about 83 minutes to train, and the embedding time took about 98% of this amount of time. Openia’s inference process is extremely fast, requiring only 0.6 milliseconds per generated code unit. This highlights the feasibility of Openia for real-time, in-process assessment of code correctness during LLM generation, making it well-suited for practical deployment scenarios such as in interactive and real-time coding environments.
@@ 5.5 Threats to Validity
Internal Validity. Threats to internal validity include the hyperparameter selection for both Openia and the baselines. To mitigate this threat, we systematically explored various hyperparameter configurations for Openia and evaluated its performance under varied conditions (as detailed in Sec. 5). For baseline methods, we reused implementations and settings from the original papers, such as those for CodeBERT and CodeT5+ [44, 46], ensuring fair comparison. Additionally, potential threats could arise from the procedures used to extract the internal representations of LLMs and their correlation with code correctness. To address this, we utilized the widely used Pytorch libraries to extract and interpret the intermediate states of LLMs during code generation. These methods were validated on diverse benchmarks and cross-verified using established practices in the field. To further minimize the risk of errors in implementation, our codebase has undergone extensive internal review, and we have made it publicly available [36].
@@
Construct Validity. Threats to construct validity relate to the appropriateness of our evaluation metrics and procedures. We employed widely recognized metrics, including Accuracy, Precision, Recall, and F1-Score, to measure the performance of the approaches, as these are standard benchmarks in evaluating classification models. Additionally, we designed experiments to include both standalone and repository-level code generation tasks, ensuring comprehensive evaluations. However, our evaluation primarily relies on automated metrics and does not yet incorporate human assessments. While automated metrics provide objectivity and scalability, human evaluation could offer deeper insights into the practical utility of Openia in real-world coding scenarios. We plan to include such assessments in future work, engaging experienced developers to evaluate the generated code based on factors like functionality, readability, and adherence to programming standards. Also, a potential threat is our procedure labeling the correctness of code generated by the studied code LLMs, particularly the “correct/passed” label, due to the potential low-quality tests. To mitigate this threat, we utilized the widely-used benchmarks [33, 34, 35] with adequate test sets. Another potential limitation is that our controlled experimental setup may not fully reflect real-world software development environments. To address this, we evaluated Openia on multiple representative benchmarks, covering diverse programming constructs and code generation requirements.
@@
External Validity. Threats to external validity concern the generalizability of our results to different LLMs and application domains. To mitigate this, we selected three widely used and representative open-source code LLMs: DeepSeek Coder, Code Llama, and Magicoder [16, 17, 18]. These models have demonstrated strong performance across diverse code-related tasks, enhancing the applicability of our findings within the code generation domain. A further potential threat arises from our focus primarily on models with parameter sizes up to 7B, due to hardware constraints. While this ensures feasibility within our experimental setup, it may limit the applicability of our conclusions to larger-scale models. As part of future work, we plan to investigate the scalability of Openia to larger LLMs and expand our experiments to a broader range of programming languages and codebases to enhance the generalizability of our approach. One potential threat to the robustness of Openia lies in the variability of input prompts. However, this threat can be partially mitigated by our evaluation across multiple benchmarks, i.e., HumanEval, MBPP, and DevEval, which exhibits diverse prompt styles and task descriptions. The consistent performance of Openia across these diverse benchmarks indicates a certain degree of robustness to prompt variation. Nonetheless, it remains important to systematically evaluate prompt sensitivity and we plan to investigate its impact on Openia’s performance in future work.